### v2

# Approximate quantum error correcting codes from conformal field theory

Abstract
The low-energy subspace of a conformal field theory (CFT) can serve as a quantum error correcting code, with important consequences in holography and quantum gravity. We consider generic 1+1D CFT codes under extensive local dephasing channels and analyze their error correctability in the thermodynamic limit. We show that (i) there is a finite decoding threshold if and only if the minimal nonzero scaling dimension in the fusion algebra generated by the jump operator of the channel is larger than $1/2$ and (ii) the number of protected logical qubits $k\geq\Omega(\log\log n)$, where $n$ is the number of physical qubits. As an application, we show that the one-dimensional quantum critical Ising model has a finite threshold for certain types of dephasing noise. Our general results also imply that a CFT code with continuous symmetry saturates a bound on the recovery fidelity for covariant codes.
Introduction.– Quantum information is fragile and can be lost when subject to decoherence. In order to robustly store and manipulate quantum information against noise, one embeds the logical qubits into a larger set of physical qubits, forming quantum error correcting codes (QECC) [1, 2]. QECCs provide promising routes to fault-tolerant quantum computing [3, 4] and have been recently realized on quantum simulators [5]. On the theoretical side, code properties of QECCs provide insight into patterns of long-range quantum many-body entanglement, connecting quantum information with condensed matter physics [6, 7] and quantum gravity [8, 9]. More recently, studies have suggested a close relationship between mixed-state topological phases and the ability of the QECCs to correct certain errors [10, 11]. It is thus important, both from a practical and a theoretical point of view, to understand which physical systems can serve as QECCs and which errors are correctable given the system or code.
Most previous studies focus on stabilizer codes [2]. The stabilizer formalism offers a natural way for decoding, where in some cases the optimal decoder can be found [6]. However, the notion of QECC goes far beyond stabilizer codes. In order to define a QECC, one may specify a $D$-dimensional code subspace spanned by a set of mutually orthogonal codeword states, $\{|\phi_{\alpha}\rangle,~{}\alpha=1,2,\cdots,D\}$, where $D=2^{k}$ to encode $k$ logical qubits. In the seminal work [12], the codeword states are chosen as the eigenstates of a local Hamiltonian. The code properties can then be related to physical properties of the system in an energy window. Another example that goes beyond stabilizers is AdS/CFT correspondence, where the code subspace is the CFT low-energy subspace corresponding to bulk graviton excitations. For this code, the code properties reveal important aspects of quantum gravity [13]. Going beyond stabilizer codes poses significant challenges to determine error correctability as there are usually no explicit decoders. We note that there are existing information-theoretic criteria for approximate error correctability [14, 15, 16] which have been used in a variety of contexts.
In this work, we show that generic CFTs furnish approximate quantum error correcting codes [17, 18], where the code subspace is given by the low-energy subspace of the CFT. As opposed to the holographic correspondence, here the CFT can be realized by a simple spin chain at criticality, such as the transverse field Ising model. We analyze the decodability of CFT codes under a finite-time evolution of a translation-invariant Lindbladian, such as uniform dephasing noise. Our main result is that the errors can be corrected at a finite threshold if and only if $\Delta_{\min}>1/2$, where $\Delta_{\min}$ is the minimal nonzero scaling dimension of CFT operators in fusion algebra generated by the Lindbladian jump operator. A natural corollary is that a CFT code can correct uniform depolarization noise at a finite threshold if and only if the algebra generated by the jump operator only contains CFT operators with scaling dimension larger than $1/2$. In addition, we show that for any CFT code the number of protected logical qubits $k\geq\Omega(\log\log n)$, where $n$ is the number of physical qubits. As an example, we analyze the Ising CFT code realized by the low-energy subspace of the critical transverse-field Ising model (TFIM) in one dimension. We show that for dephasing noise, the code does not have a finite threshold for $X$ dephasing but can correct $Y$ and $Z$ dephasing up to maximal strength.
Our work is in part inspired by recent developments in decoherence-induced mixed-state phases [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40]. While previous approaches mainly focus on finite number of replicas of the system (for computing Renyi entanglement quantities), our work directly addresses the replica limit, which is provably related to the decoding transition. As a byproduct, this work also features an example where the replica limit has fundamentally different physics from other integer Renyi indices. In contrast to previous works on QECCs in gapless systems which mainly focus on code distances [41, 42, 12, 14], our results involve the decoding threshold for extensive noise. Furthermore, our results are applicable to generic CFTs as opposed to specific models.
CFT code– Let us start with properly defining the CFT code. Given a critical quantum spin chain with $n$ spins and Hamiltonian $H$, the low-energy physics is described by a CFT. Each low-energy eigenstate $|\phi_{\alpha}\rangle$ corresponds to a scaling operator $\phi_{\alpha}$ in the CFT due to the state-operator correspondence. The energy of the state is given by $E_{\alpha}=\frac{2\pi}{n}(\Delta_{\alpha}-c/12)$, where $\Delta_{\alpha}$ is the scaling dimension and $c$ is the central charge. As $n\rightarrow\infty$, there are infinite number of low-energy eigenstates whose energies are degenerate. The encoding circuit can be chosen as a MERA tensor network [43, 44, 45, 46, 42], where the logical qubits are located at top layers.
In order to discuss the error-correcting properties, we specify the error model to be a finite-time evolution $e^{t\mathcal{L}}$ of a local Lindbladian $\mathcal{L}$, where $\mathcal{L}(\rho)=\sum_{i}\left[L_{i}\rho L^{\dagger}_{i}-\frac{1}{2}\{L^{\dagger}_{i}L_{i},\rho\}\right]$ and $L_{i}$ are local jump operators. More concretely, we consider two types of noise, dephasing and flagged dephasing. For both error models, the noise channel $\mathcal{N}=\otimes_{j}\mathcal{N}^{[j]}$ is a product of local channels $\mathcal{N}^{[j]}$ acting on a single spin $j$. A dephasing channel of strength $p$ is defined by
$\mathcal{N}^{[j]}_{p,\alpha}(\rho)=\left(1-\frac{p}{2}\right)\rho+\frac{p}{2}\sigma^{[j]}_{\alpha}\rho\sigma^{[j]}_{\alpha},$ (1)
where $\sigma^{[j]}_{\alpha}$ is the Pauli operator acting on site $j$ and $\alpha=x,y,z$. In terms of Lindbladian, this corresponds to $n$ jump operators $L_{i}=\sigma^{[i]}_{\alpha}/\sqrt{2}$ and evolution time $t=-\log(1-p)$. A flagged dephasing channel (also known as heralded noise in quantum optics [47]) of strength $p$ is defined by
$\mathcal{N}^{[j]}_{p,\alpha;\mathrm{F}}(\rho)=\left(1-p\right)\rho\otimes|0_{F}\rangle\langle 0_{F}|+p\mathcal{N}^{[j]}_{1,\alpha}(\rho)\otimes|1_{F}\rangle\langle 1_{F}|,$ (2)
where we have introduced a “flag” qubit $F$ at each site $j$. The flagged dephasing means that each site has a probability $p$ to undergo a complete dephasing, but we know which sites are subject to error by measuring the flag qubit. The flagged noise has $2n$ jump operators, $L_{2i-1}=I\otimes\sigma^{[i]}_{+}/\sqrt{2},L_{2i}=\sigma^{[i]}_{\alpha}\otimes\sigma^{[i]}_{+}/\sqrt{2}$.
We say the noise is correctable if there exists a decoding channel $\mathcal{D}$ at each size $n$ such that $\lim_{n\rightarrow\infty}F(\mathcal{D}\circ\mathcal{N}(\rho),\rho)=1$ for any $\rho$ in $\mathcal{H}_{\mathrm{code}}$, where $F$ is the Uhlmann fidelity. The uniform dephasing is harder to correct than the flagged dephasing, as one can discard the flag to obtain the uniformly dephased state, $\textrm{Tr}_{\mathrm{F}}\mathcal{N}_{p,\alpha;\mathrm{F}}(\rho)=\mathcal{N}_{p,\alpha}(\rho)$.
Error correcting condition.– In order to characterize the recoverability of a QEC, Ref. [48] introduces the entanglement fidelity for a decoding channel $\mathcal{D}$, $F_{e}=\langle\psi_{RQ}|\mathcal{D}\circ\mathcal{N}(|\psi_{RQ}\rangle\langle\psi_{RQ}|)|\psi_{RQ}\rangle$, where
$|\psi_{RQ}\rangle=\frac{1}{\sqrt{D}}\sum_{\alpha=1}^{D}|\alpha\rangle_{R}|\phi_{\alpha}\rangle_{Q}$ (3)
is the maximal-entangled state between the reference $R$ and the code subspace of the physical qubits $Q$. The fidelity between any logical state and the recovered state is lower-bounded by $F_{e}$. Let $\rho_{RQ}=\mathcal{N}(|\psi_{RQ}\rangle\langle\psi_{RQ}|)$ (see Fig. 1), one can further define the coherent information
$I_{c}=S_{Q}-S_{RQ},$ (4)
where $S_{A}:=-\textrm{Tr}(\rho_{A}\log\rho_{A})$. It has been shown that if the coherent information is close to its maximal value, that is, $I_{c}=\log D-\epsilon$, then there exists a decoding channel $\mathcal{D}$ which achieves entanglement fidelity $F_{e}>1-2\sqrt{\epsilon}$  [^1] There is also an upper bound of $F_{e}$ by $1-F_{e}>O(\epsilon)$ due to monotonicity of coherent information and Fannes inequality.. A QEC can perfectly correct the error if and only if $I_{c}=\log D$. If $\epsilon>0$ then the code is an approximate QEC.
Figure 1: (Left) The coherent information of the CFT code is defined by $I_{c}=S_{Q}-S_{RQ}$ for the state $\rho_{RQ}$ in the figure, where $W_{\mathrm{encode}}$ is an encoding isometry which outputs states in the CFT low-energy subspace and $\mathcal{N}$ denote noise channels acting on all physical sites. (Right) Coherent information of correctable (blue) and uncorrectable (red) noise for the CFT code. The correctability condition is $\Delta_{\min}>1/2$, where $\Delta_{\min}$ is the smallest scaling dimension in the fusion algebra of the noise channel jump operators.
A CFT code can approximately correct any noise that only acts on one lattice site. The reason is that the jump operator $L_{i}$ on site $i$ connects different codeword states with an amplitude [50, 51]
$\langle\phi_{\beta}|L_{i}|\phi_{\alpha}\rangle=\left(\frac{2\pi}{n}\right)^{\Delta}C_{\alpha\beta L},$ (5)
where $\Delta$ is the scaling dimension of the operator $L_{i}$ and $C_{\alpha\beta L}$ is the operator product expansion (OPE) coefficient that is independent of $n$. This violation of Knill-Laflamme condition implies that the best entanglement infidelity decreases to zero polynomially in $n$ [14].
Below we consider the CFT code under noise that uniformly acting on all sites, which is much harder to treat than the single-site noise. The coherent information $I_{c}(p,n)$ depends on the noise rate and the system size. The noise is correctable at error rate $p$ if and only if
$\lim_{n\rightarrow\infty}I_{c}(p,n)=\log D~{}(\mathrm{correctable~{}condition}).$ (6)
Given a noise model, such as flagged dephasing Eq. (2) or unflagged dephasing Eq. (1), the code is said to have a threshold $p_{c}$ if all noise channels with $p<p_{c}$ are correctable.
Perturbative expansion of coherent information– In order to derive our main result, we make use of a perturbative expansion of coherent information together with the scaling hypothesis. The derivation is sketched below.
The crucial assumpution is the scaling hypothesis, that is, the coherent information has a scaling collapse of the form
$I_{c}(p,n)=f(pn^{\nu})$ (7)
as $n\rightarrow\infty$ and $p\rightarrow 0$, where $\nu$ is a constant that is analogous to the critical exponent. The same scaling form, in particular, the same exponent $\nu$, is assumed to hold for any scaling of $p$ with respect to $n$ as long as $p\rightarrow 0$. The underlying physical intuition is to treat the noisy channel as a perturbation, which flows towards or flows away from the no-noise fixed point under the renormalization group flow. Furthermore, the scaling form is also confirmed numerically, as we demonstrate later. The scaling function $f$ is monotonically decreasing and satisfies $f(0)=\log D$ and $f(\infty)<\log D$. For dephasing noise one can easily show $f(\infty)=0$. Therefore the sign of $\nu$ determines the error correctability in the thermodynamic limit. If $\nu>0$, then $I_{c}(p,n)\rightarrow 0$ in the thermodynamic limit for arbitrary $p>0$, thus $p_{c}=0$. If $\nu<0$, then the correctability condition Eq. (6) is satisfied for small enough $p$ and thus the error has a finite threshold.
Next, we perform a perturbation theory on the coherent information assuming that $p=o(1/n)$. The quantum channel can be approximated by
$\mathcal{N}_{p}(\rho)=\rho+\frac{p}{2}\mathcal{L}(\rho)+O(p^{2}),$ (8)
where $\mathcal{L}$ is the Lindbladian. The state $\rho_{RQ}(p)=\mathcal{N}(|\psi_{RQ}\rangle\langle\psi_{RQ}|)$ can be expanded to first order in $p$. Expanding $S_{Q}$ and $S_{QR}$ to the first order (see details in appendix), we obtain
$I_{c}(p,n)=\log D+b(n)p\log p+O(p),$ (9)
where
$b(n)=\frac{1}{D^{2}}\sum_{i=1}^{n}[D\textrm{Tr}(L_{i}PL^{\dagger}_{i}P)-\textrm{Tr}(L_{i}P)\textrm{Tr}(L^{\dagger}_{i}P)]$ (10)
and $P=\sum_{\alpha}|\phi_{\alpha}\rangle\langle\phi_{\alpha}|$ is the projector onto the code subspace. The noise is correctable only if $\lim_{n\rightarrow\infty}b(n)=0$. It is worth noting that $b(n)\geq 0$ and the equality holds if and only if $PL_{i}P\propto P$, which is part of the Knill-Laflamme conditions for a single-qubit error. Higher-order perturbations can reproduce Knill-Laflamme conditions with multi-qubit errors.
Finally, using Eq. (5) we obtain $b(n)\propto n^{1-2\Delta}$, where $\Delta$ is the scaling dimension of the jump operator. The only way that $b(n)\propto n^{1-2\Delta}$ can be compatible with the scaling form $I_{c}(p,n)=f(pn^{\nu})$ is that
$\nu=1-2\Delta.$
Crucially, the scaling hypothesis allows us to infer the exponent $\nu$ for $p=O(1)$ from finite-order of perturbation theory performed at $p=o(1/n)$. Thus the error correctability condition $\nu>0$ holds at finite $p$ if and only if $\Delta>1/2$.
Expanding to higher orders of perturbation theory, we find that correctability requires that all the OPE between $L_{i}$ and itself only contains operators with scaling dimension larger than $1/2$, excluding the identity operator. If the lowest operator with scaling dimension $\Delta_{\min}$ appears at fusion order $r$, then $\nu=(1-2\Delta_{\min})/r$. We provide a calculation to the second order and an argument for higher orders in the appendix.
In order to verify our result, we numerically compute coherent information for both unflagged (Eq. (1)) and flagged (Eq. (2)) dephasing noise. We remark that two noise models should have the same $\nu$ according to our argument, because they contain the same jump operators acting on the physical system. For the unflagged case, we are restricted to small system sizes due to exponentially growing simulation complexity. For the flagged case, we can access much larger system sizes by sampling over measurement trajectories, which is a trick first explored in Ref. [52] and detailed in appendix.
Figure 2: Coherent information of the Ising CFT code under flagged and flagged dephasing with $p\leq 0.2$
Example: Ising CFT code.– As an example, let us consider the Ising CFT realized by the transverse field Ising model $H=-\sum_{i}(\sigma^{[i]}_{x}\sigma^{[i]}_{x}+g\sigma^{[i]}_{z})$ at $g=g_{c}=1$. It has three primary operators labelled by $I,\sigma,\varepsilon$ with scaling dimensions $\Delta_{I}=0,\Delta_{\sigma}=1/8,\Delta_{\varepsilon}=1$ [53]. The code subspace can be minimally chosen to be spanned by the ground state $|I\rangle$ and the second excited state $|\varepsilon\rangle$. The three single-site Pauli operators $\sigma_{x},\sigma_{y},\sigma_{z}$ corresponds to $\sigma,\partial\sigma,\varepsilon$ in the CFT [51]. We can read off the exponent $\nu$ for the $X$, and $Z$ dephasing $\nu_{x}=1-2\Delta_{\sigma}=0.75$ and $\nu_{z}=1-2\Delta_{\varepsilon}=-1$. For $Y$ dephasing, we can use second order perturbation theory to show that $\nu_{y}=(1-2\Delta_{\varepsilon})/2=-0.5$. The underlying reason is that the OPE $\partial\sigma\times\partial\sigma$ contains an operator $\varepsilon$ with $\Delta_{\varepsilon}<\Delta_{\partial\sigma}$, and further fusion does not produce a lower one than $\varepsilon$. Thus our theory predicts that the $Y$ and $Z$ dephasing have a finite threshold and $X$ dephasing has zero threshold.
We compute the coherent information for the unflagged dephasing up to $n=16$  [^2] We use MPO to store the purified state $|\psi\rangle_{RQE}$ and coarse-grain $Q$ and $E$ separately. The effective physical dimension must increase exponentially with subsystem sizes, so we have limited system size $n\leq 16$. and flagged dephasing up to $n=128$  [^3] For the sampling algorithm we use 56000 samples. with $p\leq 0.2$, see Fig. 2. The scaling collapses fit best with $\nu_{x}\approx 0.65,\nu_{y}\approx-0.5,\nu_{z}\approx-1.0$, which are close to the theoretical values.
In order to obtain the threshold, we compute $I_{c}$ for flagged noise at large $p$ and find that the threshold $p=1$ for both $Y$ and $Z$ dephasing. In Fig. 3 we show the extrapolation of $I_{c}$ to the thermodynamic limit and show that $I_{c}=\log 2$ up to $p\leq 0.8$. We also perform a scaling collapse near $p=1$ to confirm that it is indeed an unstable RG fixed point, which indicates that $p=1$ is the threshold. For unflagged dephasing, the threshold can be estimated by the crossing point of the curves $I_{c}(p)$ of different $n$. Our small-size simulation suggests that the threshold is also $p_{y}=p_{z}=1$, see appendix for details.
Figure 3: Coherent information of the Ising CFT code under $Z$ (left) and $Y$ (right) flagged dephasing with $0.2\leq p\leq 0.8$
We see that the correctable and uncorrectable dephasing of the Ising CFT code is identical to the repetition code. It is not unexpected since the repetition code can be realized by the spontaneous symmetry broken phase $10<g<$ 1 of TFIM. However, as opposed to the repitition code where the number of protected logical qubits is $k=O(1)$, the CFT code can protect more logical information as we increase $n$, as we demonstrate below.
Dimension of logical subspace.– The low-energy subspace is infinite-dimensional as $n\rightarrow\infty$, as each primary operator has infinitely many descendants. It is then natural to ask how large the code subspace can be chosen to maintain a finite threshold for correctable dephasing noise with $\Delta>1/2$. Firstly, as $b(n)\propto n^{1-2\Delta}$ in Eq. (10) for any finite-dimensional code subspace, we can store at least constant number of qubits. Below we further argue that the dimension of code subspace can scale at least as $\mathrm{polylog}(n)$, thus $k\geq\Omega(\log\log n)$. Let us choose the code subspace to be spanned by a primary state $|\phi\rangle$ and its global descendants $|\partial^{m}\bar{\partial}^{\bar{m}}\phi\rangle$, where $m$ and $\bar{m}$ are natural numbers smaller than a cutoff $M$. Then the projector to the code subspace is
$P=\sum_{m,\bar{m}=0}^{M-1}|\partial^{m}\bar{\partial}^{\bar{m}}\phi\rangle\langle\partial^{m}\bar{\partial}^{\bar{m}}\phi|$ (11)
As we show in the appendix, if $M=\mathrm{polylog}(n)$, then $\lim_{n\rightarrow\infty}b(n)\rightarrow 0$ given any $\Delta_{K}>1/2$. The total dimension of code subspace $2^{k}=M^{2}$, which means $k\propto\log(M)=O(\log\log n)$. Thus, we establish that the number of logical qubits increases with the number of physical qubits. In principle one can encode more logical qubits by including Virasoro descendants in the code subspace. By including all descendants below scaling dimension $M=\mathrm{polylog}(n)$, one may reach $k=O(\mathrm{polylog}(n))$ by the Cardy formula [50]. We leave it to future work to investigate its error protection properties.
Covariant CFT code.– Let us consider a particular case of the CFT code where the CFT has a global $U(1)$ symmetry. Then there exists a conserved charge density which has scaling dimension $\Delta=1$. The dephasing noise on the conserved charge is a correctable error since $\nu=1-2\Delta=-1$. The coherent information gives $1-F_{e}>\sqrt{\delta I_{c}}\propto\sqrt{p}n^{-1/2}$ at small $p$. The same conclusion holds for arbitrary dimensions. On the other hand, it has been shown in Refs. [56, 57] that $1-F_{e}\leq O(n^{-1/2})$ for any covariant code with $U(1)$ symmetry. Thus, the CFT code with $U(1)$ symmetry saturates the covariant code bound.
Discussion.– Our result can be readily applied to depolarization noise, where all Pauli errors can happen. The CFT code can correct depolarization noise at a finite threshold if and only if all scaling dimensions $\Delta>1/2$. Such a CFT does not exist for Virasoro minimal models and all $c=1$ free boson models. CFTs that satisfy this condition have $c>1$ and may only be realized beyond spin-$1/2$ chains and $2-$local interactions [58]. One such example is the monster CFT with the lowest primary $\Delta=4$, which has been conjectured to be dual to a bulk quantum gravity theory [59]. Another example which has $\Delta_{\min}>1/2$ is the $SU(N)_{1}$ WZW model for $N\geq 3$, where the simplest case $N=3$ can be realized by a spin-1 chain [60]. One interesting feature of these CFT codes is that they may have a finite threshold in one spatial dimension, which cannot be achieved by the ground state subspace of any gapped local Hamiltonian.
In terms of characterization of mixed-state phases, one of our technical contributions is the perturbative calculation of coherent information without applying replica trick, while prior works mostly focus on finite replica calculations. In fact, the Renyi coherent information may give incorrect predictions of the code properties. For example, it gives a wrong decoding threshold for the toric code [20]. This work has given another example which shows that working with integer Renyi index predicts wrong error-correctability of infinitesimally small noise strength. For the Ising CFT code, uniform $Z$ dephasing can be corrected at a finite threshold but the Renyi coherent information drops as long as $p\neq 0$, see appendix for details.
Our work opens up several future directions. Firstly, even though we have shown that some errors can be corrected for CFT code, it is still nontrivial to construct an explicit decoder, an important ingredient of QECC. Secondly, we can consider logical operations for CFT code and see if it can be performed fault-tolerantly. Thirdly, it is interesting to consider the bulk dual in AdS/CFT with boundary dephasing [61, 62], which may inspire the solution of the two questions above.
Acknowledgements.– We thank Charles Cao, Yu-Hsueh Chen, Tarun Grover, Yangrui Hu, Isaac Kim, Yaodong Li, David Lin, Peter Lu, Ruochen Ma, Alexey Milekhin, Daniel Parker, Bowen Shi, Guifre Vidal, Sisi Zhou, Tianci Zhou and Zheng Zhou for helpful discussions. Y.Z. in particular thanks Patrick Hayden for his insight that leads this work. This work was supported by the Perimeter Institute for Theoretical Physics (PI) and the Natural Sciences and Engineering Research Council of Canada (NSERC). Research at PI is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development Canada and by the Province of Ontario through the Ministry of Colleges and Universities. This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP).